In a ∆ABC, if A = 4π/9 and B = 2π/9 , then a² - b² is equal to please solution

Grade:11

2 Answers

Rajan Sahu
6 years ago
Here a / sin 80°= b / sin 40°(Applying sine rule) Now a/2cos40°=bSo cos 40° = a/2b.. Now applying cosine rule=>Cos 40°=( a^2+c^2-b^2) /2acSo a/2b= (a^2 + c^2 - b^2)/2ac=> a^2 c = a^2 b + c^2 b - b^3=>a^2(b-c) = b(b ^2 - c ^2)=>a^2 = b ( b+c) =>a^2 - b^2 = bc
Joynto
6 years ago
A=4pie/9 & B=2pie/9by applying a square - b square formula we get16pie square/81 -4pie square/81=12pie square/81=4pie square/81 [ dividing 3 by numerator & denominator]